# There is No Silver Bullet: Benchmarking Methods in Predictive Combinatorial Optimization

### v4

@@ 4.5 Key factors for future PnO methodology design
RQ1: Relationships of prediction accuracy and decision quality In common sense, a better prediction would lead to a better subsequent decision. Besides a case study shown in [18], we explore the prediction loss and decision objective in both PtO and PnO. We illustrate the learning curve on the real-world knapsack (energy) dataset by plotting the prediction loss (Mean-squared error) and evaluation results (regret) of one method for each category, shown in Figure 5(c d). We visually scale the prediction loss of Blackbox to $5*10^{9}$ for clarity, while others to $10^{8}$.
It is observed that PnO training methods like SPO, listwise-LTR, and LODL achieve lower regret than the vanilla PtO method. The prediction loss of LODL shares a similar pattern with PtO, while LTR’s is different. This demonstrates despite higher prediction loss incurred in the end-to-end training, these PnO methods obtain better decisions by their own loss functions concerning the ultimate decision objectives. This might correspond to the proposition that PnO achieves better decision quality by finding better error trade-offs [10] by information of decision objectives.
@@
RQ3: Versatility of PnO under various parameters We explore the versatility of PnO as follows. (1) Figure 6 examines the impact of varying constraint on the training of KG dataset. We set the capacity of 30, 60, and 90. It was observed that as the constraints became more restrictive, the relative regrets became higher. This observation potentially indicates that constraint satisfaction is a prevalent factor constraining the performance of current PnO training methods. (2) Figure 7(V1~V2) show the performance when the number of decision variables increases to 40 and 60. With increasing variable size, the two-stage approach remains stable under different problem settings, though its final decision quality is inferior to some of the PnO training models. Conversely, some PnO models exhibited fluctuations and degradation, such as discrete-category methods and point-wise LTR. (3) Figure 7(G1~G2) show the generalizability when the PnO models are trained on one problem setting (e.g., constraint of capacity 30) and tested on the other one (e.g., constraint of capacity 60 or 90). Similar performance degradation occurs in some PnO models, including discrete-category and statistical-category models, in this setting. This indicates current PnO approaches may perform inadequately when facing more stringent constraints, larger decision variable sizes, and the need for generalization across diverse optimization parameters. More detailed results are shown in Appendix C.2.
@@ 5 Conclusion
Appendix A Model Details
A.1 Prediction Model
@@
where $\mathbf{a}^{(1)}=\mathbf{x}$ and $\mathbf{y}=\mathbf{a}^{(K)}$ are input and output for $\mathcal{M}$, $a^{i}$ is the hidden vector for $i=2,\cdots,K-1$, $\mathbf{W}$ is the weight term, $b$ is the bias term and $\sigma$ is the activation function where we adopt ReLU. In experiments, we adopt $K=3$ layers, and the size of intermediate hidden units is set as 32.
A.2 Two-stage approach in PtO
The two-stage approach is the vanilla model used when the coefficients of the optimization problem are uncertain and require prediction. In the prediction stage, the predictor estimates the coefficients of the optimization problem, and optimizes with the original prediction loss using pre-collected labeled data pairs $D=\{x_{i},y_{i}\}$. In the optimization stage, the coefficient on the test data is first inferred so that the optimization problem becomes certain. The vanilla solver is used to solve the optimization problem.
@@ A.2 Two-stage approach in PtO
A.3 Details for PnO models
A.3.1 The Discrete category
@@ A.3.4 The Surrogate Category
Appendix B Problem and Dataset Details
For most problems, we adopt generated data or data that has been publicly available. For our new benchmark combinatorial advertising problem, which will be released, we collect data that contain encrypted and processed personal features and mobile application activity records, and all the used data are permitted by the users. In the following, we elaborate on the detailed data processing procedure for each problem. We also give code framework 1 with base classes. The full code and processed dataset will be released after publication.
B.1 Energy-cost aware Scheduling
@@ B.6 Portfolio Optimization
B.7 Combinatorial Advertising for Inclusive Finance
@@
B.7.1 Predict-and-optimize formulation
@@
B.7.2 Data processing and model details
@@
$\bullet$    Optimization: For the Integer Linear Programming formulation, we adopt OR-Tools solver [43] with SCIP (Strategic Consortium of Intelligence Professionals) [8, 9] backend has been utilized for the resolution of the discrete linear programming problem. Specific parameter values have been assigned to the optimization process, including channel costs (0.5 for channel 1, 1 for channel 2), costs for the combined strategy [0, 0.5, 1.0, 1.5] (which is simply treated as the sum of channel costs), and budget constraints, where the total budget is derived from the product of the number of users and the per-capital budget allocation of 0.1.
B.7.3 Evaluation
@@
where $w$ indicates the treatment and $y$ indicates the label. In our case, the treatment group means the group of people whose combinatorial advertisement strategy is identical to the offline data, whereas the control group means the opposite.
B.7.4 License, Use Terms and Privacy
In the following terms, the company that provides the data for combinatorial advertising is referred to as “The Corporation”.
Data use terms and license
@@ B.7.4 License, Use Terms and Privacy
“The Corporation” strives to enhance public access to and use of data that it collects and publishes, to promote data utilization and academic research. The data are organized in datasets (the“Datasets”). The Datasets are collections of data, managed by “The Corporation” and provided in a number of machine-readable formats. “The Corporation” provides you with access to the Datasets free of charge subject to the terms of this agreement (these “Dataset Terms”). Use of data derived from the Datasets, which may appear in formats such as tables and charts, is also subject to these Dataset Terms.
Comment
We encourage you to share your suggestions and ideas for using or facilitating access to the Datasets with “The Corporation” by sending an email to xxx111111will be released later..
Licenses and Restrictions
@@
You may not use the name, any trade-mark, official mark, official emblem or logo of “The Corporation”, or any of its other means of promotion or publicity, without “The Corporation”’s prior written consent nor in any event to represent or imply an association or affiliation with “The Corporation”.
@@
“The Corporation” reserves the right at any time and from time to time to modify or discontinue, temporarily or permanently, this website, the Datasets, any means of accessing or utilizing the Datasets, or the API, at our sole discretion with or without prior notice to you.
“The Corporation” may at our sole discretion, under any circumstances, for any or no reason whatsoever and with or without prior notice to you, terminate your access to the Datasets, any means of accessing or utilizing the Datasets or the API.
“The Corporation” DISCLAIMS ALL WARRANTIES OF ANY KIND RELATED TO THE PROVISION OF THE DATASETS AND THE APIs.
@@
“The Corporation” SHALL NOT BE RESPONSIBLE OR LIABLE TO YOU FOR ANY LOSS OR DAMAGE OF ANY SORT INCURRED BY YOU IN CONNECTION WITH YOUR USE OF THE DATASETS. “The Corporation” also shall not be responsible or liable for the accuracy, usefulness or availability of any data in the Datasets.
You acknowledge that these Dataset Terms constitute a non-exclusive agreement. “The Corporation” may develop products or services that compete with products or services that you offer without incurring any liability.
Other parties may have ownership interests in some of the data and information (“Materials”)  contained on the Site. “The Corporation” in no way represents or warrants that it owns or controls all rights in all Materials, and “The Corporation” will not be liable to you for any claims brought against you by third parties in connection with your use of any Materials.
These Dataset Terms may be amended by “The Corporation” from time to at our sole discretion. If there is a Chinese version of the Dataset Terms, the Chinese Dataset Terms shall prevail where the Chinese Dataset Terms conflict with the English Terms.
By continuing to use the Datasets subsequent to “The Corporation” making available an amended version of these Dataset Terms, you acknowledge, agree and consent to such amendment.
@@
Privacy
Safeguarding user information and protecting user privacy is paramount in “The Corporation”’s operation since its inception. The Company has established a comprehensive administrative mechanism and standardized employee training system for stringent information security management. “The Corporation” has also been deploying innovative technologies to promote user data protection. For example, the Company launched a Smart Finance Institute in 2018 for research and development in the field of artificial intelligence that can be applied in various aspects of financial services. In addition, “The Corporation” is also a member of the National Information Security Standardization Technical Committee and Mobile Application (APP) Security Committee, maintaining up to date knowledge and compliant regarding the latest cyber-security regulatory requirements.
Appendix C Experimental Details
C.1 Analysis on Benchmark Results
@@ C.2 Details of versatility of PnO in RQ3
C.3 Sensitivity analysis of training on hyperparameters
@@
C.4 Theoretical Interpretation on benchmarks
@@
Appendix D Terms of Use, Limitations and Broader Impacts
@@
As part of our benchmark work, we utilized numerous datasets in the benchmarking process, and we adhere to the respective usage terms and licenses elaborated in Appendix B. Additionally, we are releasing a new industrial dataset, and the license, use terms, and privacy are also elaborated in Appendix B.7.4. We affiliate a code framework in supplemental materials and full code and data will be made public after publication.
@@
The limitations of this paper include the inherent difficulty in analyzing the internal mechanisms of PnO. Aside from a few existing works that examine the theoretical underpinnings of PnO (like SPO [18]), at present, it is challenging to provide theoretical explanations or straightforward criteria to determine when PnO outperforms PtO or which specific model is the best choice. Therefore, our focus for now is primarily on the empirical results of these methods on current datasets. To assist practitioners, we have categorized and analyzed these methods based on their generalizability, usage conditions, and associated costs, and our experimental results. In Appendix C.4, we also attempt to interpret the benchmark results based on existing theories for the reader’s reference. We offer a predict-and-optimize model using and deployment recommendations in Sec. 3.5 based on these above considerations for practitioners.
@@
As far as we are concerned, this paper does not have critical negative societal impacts. We hope that this benchmark work can support better decision-making under uncertainty by utilizing end-to-end prediction and optimization methods for training predictive models more appropriately. In Appendix B, we also discuss the possible positive implications of these decisions, including more energy-efficient production scheduling models, more rational resource allocation under uncertain scenarios, more appropriate social matching, more interpretable models, more appropriate capital allocation, and more inclusive financial services. Certainly, we acknowledge that similar to previous works on predict-and-optimize (PnO), decision-making under uncertainty could potentially be applied to inappropriate purposes. However, given the current challenges of deploying PnO in industry and practice, we consider the negative impacts will be minimal. We hope that this benchmark can facilitate future advances in this field and aid in creating better decisions under uncertain scenarios for people.

### v5

@@ 4.5 Key Factors for PnO Methodology Design
RQ1: Relationships of prediction accuracy and decision quality In common sense, a better prediction would lead to a better subsequent decision. Besides a case study shown in [20], we explore the prediction loss and decision objective in both PtO and PnO. We illustrate the learning curve on the real-world knapsack (energy) dataset by plotting the prediction loss (Mean-squared error) and evaluation results (regret) of one method for each category, shown in Figure 5(c d).
It is observed that PnO training methods like SPO, listwise-LTR, and LODL achieve lower regret than the vanilla PtO method. The prediction loss of LODL shares a similar pattern with PtO, while LTR’s is different. This demonstrates despite higher prediction loss incurred in the end-to-end training, these PnO methods obtain better decisions by their own loss functions concerning the ultimate decision objectives. This might correspond to the proposition [9] that PnO achieves better decision quality by finding better error trade-offs by information of decision objectives. We show more qualitative analysis in Appendix D.5.
@@
RQ3: Versatility of PnO under various parameters We explore the versatility of PnO as follows. (1) Figure 6 examines the impact of varying constraint on the training of KG dataset. We set the capacity of 30, 60, and 90. It was observed that as the constraints became more restrictive, the relative regrets became higher. This observation potentially indicates that constraint satisfaction is a prevalent factor constraining the performance of current PnO training methods. (2) Figure 7(V1~V2) show the performance when the number of decision variables increases to 40 and 60. With increasing variable size, the two-stage approach remains stable under different problem settings, though its final decision quality is inferior to some of the PnO training models. Conversely, some PnO models exhibited fluctuations and degradation, such as discrete-category methods and point-wise LTR. (3) Figure 7(G1~G2) show the generalizability when the PnO models are trained on one problem setting (e.g., constraint of capacity 30) and tested on the other (e.g., constraint of capacity 60 or 90). Similar performance degradation occurs in some PnO models, including discrete-category and statistical-category models, in this setting. This indicates current PnO approaches may perform inadequately when facing more stringent constraints, larger decision variable sizes, and the need for generalization across diverse optimization parameters. More details are shown in Appendix D.2.
@@ 5 Conclusion
Appendix A Related Work
Despite the significant progress neural networks have made in machine learning, there is still substantial room for improvement when it comes to solving optimization problems. Although several studies [55, 34, 63, 12, 31, 42, 43, 41, 14, 27] have explored the use of neural networks for combinatorial optimization problems in deterministic environments, especially in the constrained CO tasks [71, 69], however, many of them are incompetent to catch up heuristic algorithms (such as the LKH for the traveling salesman problem) in terms of solving efficiency and decision quality. More importantly, they are not directly applicable to the optimization tasks under uncertainty, especially for unknown coefficients. Therefore, a newer area of interest involves combining machine learning with optimization techniques to tackle problems with uncertain parameters, elaborated as follows.
[7] pioneered efforts in merging predictive and prescriptive analysis for optimization under uncertainty. A notable contribution is the SPO method [20], which introduces subgradient-based surrogate functions to replace non-differentiable regret functions in linear optimization problems, and its extension, SPO-relax [46], which applies continuous relaxation to combinatorial problems and is utilized in our experiments. It is important to note that these studies primarily focus on predictive models before optimization solvers, which are often treated as standard heuristics (such as LKH3 [30] or Dijkstra [17]) or commercial solvers (like Gurobi [28]).
Recently, there has been a growing body of work on predict-and-optimize approaches (also known as decision-focused learning [72, 44, 60]). A significant trend within this category involves using relationships between solutions to optimization problems to develop better predictive models. For instance, the NCE (Noise-Contrastive Estimation) method [49] utilizes a noise-contrastive estimation technique [29] to produce predictions that aim to achieve optimal solutions with better decision quality compared to non-optimal ones. The LTR (Learning to Rank) technique [44] explores the connection between pairwise learning to rank in NCE, leading to the development of various learning-to-rank methodologies such as pointwise rank [11], pairwise rank [35], and listwise rank [10], which aim to generate predictions reflecting the relative importance of multiple solutions.
Another recent approach involves using neural network functions as surrogates for the original objective functions. Studies such as LODL [60] and EGL [61] propose learning surrogate objective functions from sample data. LANCER [77] adopts a similar strategy, integrating surrogate function learning with optimization solving and objective function learning. SurCO [23] suggests replacing the original non-linear objective with a linear surrogate, allowing the use of existing linear solvers.
In addition, some works propose implicit differentiation to obtain derivatives for convex optimization objectives for continuous optimization problems. OptNet [3] is the first work to differentiate quadratic programs, and then CPLayer [1] and more [18, 40, 19] extends this spirit to more general convex programs. QPTL [72] proposes to differentiate through combinatorial optimization tasks by relaxing the problem to the continuous pace, and IntOpt [45] improves the interior-point method so that its backpropagation differentiable. The following works  [68, 22, 52] generate this paradigm for discrete problems.
However, some of these methods are limited to specific types of predict-and-optimize problems, such as quadratic optimization objectives [3] or convex objectives. Although methods based on solution importance [49, 44] and surrogate objective functions [60, 61, 77] do not restrict the type of optimization, they require additional information, such as multiple solutions or a large number of optimization samples [60], to train the surrogate function before complete end-to-end learning can be performed. Furthermore, the most critical issue is that most of these methods struggle with combinatorial optimization due to the difficulty of differentiating through discrete decision variables, making predict-and-optimize for CO problems particularly challenging.
Besides the works mentioned above, some works also focus on the optimization under uncertainty. However, since current PnO methods do not support arbitrary uncertainty scenarios and are limited to cases with a known closed-form optimization objective but unknown coefficients, they are currently out of the scope of our benchmark. In  [36, 78], the form of the optimization objective function is uncertain and changes with time (especially for the pickup task), and [65] focuses more on the performance of regression among combinatorial spaces when there is no distribution of test data during training, but training data generation is available. Therefore, in this work, we mainly focus on methods within the predict-and-optimize paradigm and do not include them for now.
Appendix B Model Details
B.1 Prediction Model
@@
where $\mathbf{a}^{(1)}=\mathbf{x}$ and $\mathbf{y}=\mathbf{a}^{(K)}$ are input and output for $\mathcal{M}$, $a^{i}$ is the hidden vector for $i=2,\cdots,K-1$, $\mathbf{W}$ is the weight term, $b$ is the bias term and $\sigma$ is the activation function where we adopt ReLU. In experiments, we adopt $K=3$ layers and the size of intermediate hidden units is set as 32.
B.2 Two-stage approach in PtO
The two-stage approach is the vanilla model used when the coefficients of the optimization problem are uncertain and require prediction. In the prediction stage, the predictor estimates the coefficients of the optimization problem and optimizes with the original prediction loss using pre-collected labeled data pairs $D=\{x_{i},y_{i}\}$. In the optimization stage, the coefficient on the test data is first inferred so that the optimization problem becomes certain. The vanilla solver is used to solve the optimization problem.
@@ B.2 Two-stage approach in PtO
B.3 Details for PnO models
B.3.1 The Discrete category
@@ B.3.4 The Surrogate Category
Appendix C Problem and Dataset Details
For most problems, we adopt generated data or data that has been publicly available. For our new benchmark combinatorial advertising problem, which will be released, we collect data that contain encrypted and processed personal features and mobile application activity records, and all the used data are permitted by the users. In the following, we elaborate on the detailed data processing procedure for each problem.
C.1 Energy-cost aware Scheduling
@@ C.6 Portfolio Optimization
C.7 Combinatorial Advertising for Inclusive Finance
@@
C.7.1 Predict-and-optimize formulation
@@
C.7.2 Data processing and model details
@@
$\bullet$    Optimization: For the Integer Linear Programming formulation, we adopt OR-Tools solver [53] with SCIP (Strategic Consortium of Intelligence Professionals) [8] backend has been utilized for the resolution of the discrete linear programming problem. Specific parameter values have been assigned to the optimization process, including channel costs (0.5 for channel 1, 1 for channel 2), costs for the combined strategy [0, 0.5, 1.0, 1.5] (which is simply treated as the sum of channel costs), and budget constraints, where the total budget is derived from the product of the number of users and the per-capital budget allocation of 0.1.
C.7.3 Evaluation
@@
where $w$ indicates the treatment and $y$ indicates the label. In our case, the treatment group is the group of people whose combinatorial advertisement strategy is identical to the offline data, whereas the control group is the opposite.
C.7.4 License, Use Terms and Privacy for combinatorial advertising dataset
If you use the combinatorial advertising dataset provided in the benchmark, you agree to the following terms of use from “The Finvolution Group”.
C.7.4.1 Data use terms and license
@@
The Finvolution Group strives to enhance public access to and use of data that it collects and publishes, to promote data utilization and academic research. The data are organized in datasets (the“Datasets”). The Datasets are collections of data, managed by The Finvolution Group and provided in a number of machine-readable formats. The Finvolution Group provides you with access to the Datasets free of charge subject to the terms of this agreement (these “Dataset Terms”). Use of data derived from the Datasets, which may appear in formats such as tables and charts, is also subject to these Dataset Terms.
C.7.4.2 Licenses and Restrictions
@@
You may not use the name, any trade-mark, official mark, official emblem or logo of The Finvolution Group, or any of its other means of promotion or publicity, without The Finvolution Group’s prior written consent nor in any event to represent or imply an association or affiliation with The Finvolution Group.
@@
The Finvolution Group reserves the right at any time and from time to time to modify or discontinue, temporarily or permanently, this website, the Datasets, any means of accessing or utilizing the Datasets, or the API, at our sole discretion with or without prior notice to you.
The Finvolution Group may at our sole discretion, under any circumstances, for any or no reason whatsoever and with or without prior notice to you, terminate your access to the Datasets, any means of accessing or utilizing the Datasets or the API.
THE FINVOLUTION GROUP DISCLAIMS ALL WARRANTIES OF ANY KIND RELATED TO THE PROVISION OF THE DATASETS AND THE APIs.
@@
THE FINVOLUTION GROUP SHALL NOT BE RESPONSIBLE OR LIABLE TO YOU FOR ANY LOSS OR DAMAGE OF ANY SORT INCURRED BY YOU IN CONNECTION WITH YOUR USE OF THE DATASETS. The Finvolution Group also shall not be responsible or liable for the accuracy, usefulness or availability of any data in the Datasets.
You acknowledge that these Dataset Terms constitute a non-exclusive agreement. The Finvolution Group may develop products or services that compete with products or services that you offer without incurring any liability.
Other parties may have ownership interests in some of the data and information (“Materials”)  contained on the Site. The Finvolution Group in no way represents or warrants that it owns or controls all rights in all Materials, and The Finvolution Group will not be liable to you for any claims brought against you by third parties in connection with your use of any Materials.
These Dataset Terms may be amended by The Finvolution Group from time to at our sole discretion. If there is a Chinese version of the Dataset Terms, the Chinese Dataset Terms shall prevail where the Chinese Dataset Terms conflict with the English Terms.
By continuing to use the Datasets subsequent to The Finvolution Group making available an amended version of these Dataset Terms, you acknowledge, agree and consent to such amendment.
@@
C.7.4.3 Privacy
Safeguarding user information and protecting user privacy is paramount in The Finvolution Group’s operation since its inception. The Company has established a comprehensive administrative mechanism and standardized employee training system for stringent information security management. The Finvolution Group has also been deploying innovative technologies to promote user data protection. For example, the Company launched a Smart Finance Institute in 2018 for research and development in the field of artificial intelligence that can be applied in various aspects of financial services. In addition, The Finvolution Group is also a member of the National Information Security Standardization Technical Committee and Mobile Application (APP) Security Committee, maintaining up to date knowledge and compliant regarding the latest cyber-security regulatory requirements.
Appendix D Experimental Details
D.1 Analysis on Benchmark Results
@@ D.2 Details of versatility of PnO in RQ3
D.3 Sensitivity analysis of training on hyperparameters
@@
D.4 Theoretical Interpretation on benchmarks
@@
D.5 Qualitative Analysis
In Fig 5, we visually scale the prediction loss of Blackbox to $5*10^{9}$ for clarity, while others to $10^{8}$.
Refer to caption Refer to caption
(a) Values and weights of knapsack problem (b) Predicted values for various methods
Figure 10: Qualitative analysis for knapsack problem on the 10-th test instance.
We conduct a set of qualitative analyses on the knapsack problem to explore the reasons behind the success or failure of these methods. First, Fig 10(a) visualizes a bar plot of the 10-th test instance in the knapsack (gen) problem instance, showing the values and weights. We can see that the range of predicted values by SPO is roughly the same as that of the two-stage model, but the Blackbox model, without labels, has a significant difference from the true values. In Fig 10 (b) we plot the predicted values for 20 items for three models: the two-stage model, the well-performing SPO model, and the poorly-performing Blackbox model.
We can observe that the SPO model achieved the best decision quality (with the objective value of 18, selecting items 1, 3, 6, 9, 17, and 18). The two-stage model’s decision quality was slightly lower (with the objective value of 17, selecting items 1, 3, 6, 8, 9, and 17). SPO improved decision quality by replacing item 9 with item 18, by predicting a higher predicted value to item 18, and by lowering the prediction for item 9, which is guided by information of optimization objective while the two-stage does not involve this. Although this increased the prediction error, it enhanced the decision quality.
For the Blackbox model, we observed that its predicted values were similar to SPO for some items, but due to the lack of label supervision during the prediction phase (not using coefficient labels), the prediction errors were larger, leading to poorer decision quality.
Appendix E Terms of Use, Limitations and Broader Impacts
@@
Terms of Use As part of our benchmark work, we utilized numerous datasets in the benchmarking process, and we adhere to the respective usage terms and licenses elaborated in Appendix C. Additionally, we are releasing a new industrial dataset, and the license, use terms, and privacy are also elaborated in Appendix C.7.4. If readers use our benchmark dataset, please also adhere to the corresponding user terms specified by them.
@@
Limitations The limitations of this paper include the inherent difficulty in analyzing the internal mechanisms of PnO. Aside from a few existing works that examine the theoretical underpinnings of PnO (like SPO [20]), at present, it is challenging to provide theoretical explanations or straightforward criteria to determine when PnO outperforms PtO or which specific model is the best choice. Therefore, our focus for now is primarily on the empirical results of these methods on current datasets. To assist practitioners, we have categorized and analyzed these methods based on their generalizability, usage conditions, and associated costs, and our experimental results. In Appendix D.4, we also attempt to interpret the benchmark results based on existing theories for the reader’s reference.We offer a predict-and-optimize model using deployment recommendations in Sec. 3.5 based on these above considerations for practitioners.
Our work does not aim to cover all combinatorial optimization problems, predictors, solvers, or Predict-and-Optimize (PnO) methods. Instead, we present a user-friendly framework with representative implementations for each module. Readers who wish to implement their deployments in any module can do so by replacing the corresponding block within our framework.
@@
Broader Impacts As far as we are concerned, this paper does not have critical negative societal impacts. We hope that this benchmark work can support better decision-making under uncertainty by utilizing end-to-end prediction and optimization methods for training predictive models more appropriately. In Appendix C, we also discuss the possible positive implications of these decisions, including more energy-efficient production scheduling models, more rational resource allocation under uncertain scenarios, more appropriate social matching, more interpretable models, more appropriate capital allocation, and more inclusive financial services. Certainly, we acknowledge that similar to previous works on predict-and-optimize (PnO), decision-making under uncertainty could potentially be applied to inappropriate purposes. However, given the current challenges of deploying PnO in industry and practice, we consider the negative impacts will be minimal. We hope that this benchmark can facilitate future advances in this field and aid in creating better decisions under uncertain scenarios for people.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] The main claims in the abstract and introduction accurately reflects the paper’s contributions and scope.
    2. (b)
      Did you describe the limitations of your work? [Yes] We discuss the limitations in Appendix E.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] We discuss the impacts in Appendix E.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] We read the ethics review guidelines and ensure the paper conforms to them.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] We make some attempts to explain the benchmark results with existing theories in Appendix D.4, but it is not our primary focus.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] We make some attempts to explain the benchmark results with existing theories in Appendix D.4, but it is not our primary focus.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] We specified the model details in Sec. 4.3 and Appendix B. The code will be open-sourced after publication.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We specify training details in Appendix C and D
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] We report error bars for knapsack problem in Fig 7.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We specify the used computing resource in Sec 4.3.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We cite the dataset sources in Table 2, and specified all the dataset details in Appendix C.
    2. (b)
      Did you mention the license of the assets? [Yes] We discuss the license of assets in Appendix C.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We introduce our new dataset in Appendix C.7.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] We discuss in Appendix C.7.4.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We discuss in Appendix C.7.4.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] The paper does not use crowdsourcing or conducted research with human subjects.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] The paper does not use crowdsourcing or conducted research with human subjects.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] The paper does not use crowdsourcing or conducted research with human subjects.